Parth (parthkohli):

How to solve this?\[\int xe^{6x}dx \]

OpenStudy (anonymous):

use integration by parts

OpenStudy (anonymous):

u=e^6x du=6e^6x dv=x v=x^2/2 use integration by parts now

OpenStudy (lgbasallote):

i think it's easier to use \[u = x \implies du = dx\] \[dv = e^{6x} \implies v = \frac{e^{6x}}6\]

Parth (parthkohli):

I don't really know much about integration by parts. Please explain.

Parth (parthkohli):

Wait, I'm getting it...

OpenStudy (anonymous):

check this link..

Parth (parthkohli):

\[\int udv = uv - \int vdu \]So that's why we take \(u = x\) and \(dv = e^{6x}dx\) right?

Parth (parthkohli):

Hello, Ms. Chlorophyll!

Parth (parthkohli):

What should I do next?

OpenStudy (lgbasallote):

you derive u and integrate dv

OpenStudy (anonymous):

Go with @lgbasallote's way, because what you want is to reduce the power of x

Parth (parthkohli):

Ok...\[du = dx\]Because derivative of \(x\) is 1.

OpenStudy (lgbasallote):

right.. and integral of dv?

Parth (parthkohli):

And\[\int e^{6x}dx = {e^{6x} \over 6} \]

OpenStudy (lgbasallote):

yes.

OpenStudy (lgbasallote):

so now you have \[u = x\] \[du = dx\] \[v = \frac{e^{6}}6\] you can now solve \[uv - \int vdu\]

OpenStudy (lgbasallote):

that v should be v = [e^(6x)]/6...you know what i mean

Parth (parthkohli):

So,\[\int udv = uv - \int vdu \]...then\[\int xe^{6x}dx = {xe^{6x} \over 6} - \int {e^{6x} dx} \]??

Parth (parthkohli):

I'm very slow at integration right now haha

OpenStudy (anonymous):

correct, now take integral of e^(6x) one more time!

OpenStudy (anonymous):

=xe^6x/6−e^6x/6

Parth (parthkohli):

\[\int e^{6x} = {e^{6x} \over 6}\]

Parth (parthkohli):

dx at the end...

Parth (parthkohli):

\[\int xe^{6x}dx = {xe^{6x} \over 6} - {e^{6x} \over 6} \]\[\implies {x e^{6x} - e^{6x} \over 6} \]

Parth (parthkohli):

\[\implies {e^{6x}(x - 1) \over 6} \]Is that correct?

Parth (parthkohli):

and + c

Parth (parthkohli):

Really? Is that really it!??!

Parth (parthkohli):

Wow... can't believe I did it :|

OpenStudy (anonymous):

Remember when the terms in the integral have no relation of derivative for substitute, then think about Integral by parts! Great job :)

Parth (parthkohli):

Thank you, Ms. Chlorophyll!

OpenStudy (anonymous):

Also know, choose dv as the less complicated term!

Parth (parthkohli):

OK, I get it... :)

OpenStudy (anonymous):

:) Practice makes perfect :)

Parth (parthkohli):

OK, I am gonna practice!

Parth (parthkohli):

Thank you!